OpenStudy (anonymous):

On independence day,chocolates were to be equally distributed among 360 children in a school.Actually on that day 120 children were absent,on account of this each child got 5 chocolates extra.How many chocolates each student got on that day?

OpenStudy (anonymous):

@hartnn @Hero @primeralph

OpenStudy (anonymous):

@experimentX @inkyvoyd @Preetha

hartnn (hartnn):

ok, let the number of chocolates be 'x' then if all students were present, can you tell me how may chocolates each student would get ?

OpenStudy (anonymous):

360/x

hartnn (hartnn):

wouldn't it be x/360 ?

OpenStudy (anonymous):

yeah. you are right

OpenStudy (anonymous):

then,

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

1.let the number of chocolates each student got be "x". 2.If all of them were present then each student would have got x -5 chocolates. 3.Now so total number of chocolates distributed if all were present, is = 360 * (x-5) 4.Total number of chocolates distributed when 240 children were present = 240 * x and equation 3 and 4 are equal as total chocolates distributed are same. solve this 360 * (x -5) = 240 * x and you will get the answer

OpenStudy (anonymous):

yeah. thank you